# Fermion masses and mixings and charged lepton flavor violation in a 3-3-1 model with inverse seesaw

### v1

@@ VI Lepton flavor violation decays $e_{j}\rightarrow e_{i}$
Figure 4: Correlation between the branching ratios $\text{Br}(\mu\to e\gamma)$ and $\text{Br}(\tau\to e\gamma)$ (panel (a)), and $\text{Br}(\tau\to\mu\gamma)$ and the conversion rate $\mathrm{CR}(\mu^{-}\text{Al}\to e^{-}\text{Al})$ (panel (b)). The color gradient represents the values of $\text{Br}(\tau\to\mu\gamma)$ in panel (a) and the gauge boson mass $m_{Y}$ in panel (b). The yellow shaded region corresponds to the exclusion zone from the MEG experiment [127] and the red dot dashed line indicates the current bound for $\mu\to e$ conversion in titanium (Ti) nuclei [131].
$m_{n_{4}}$ $m_{n_{5}}$ $m_{n_{6}}$ $m_{n_{7}}$ $m_{n_{8}}$ $m_{n_{9}}$ $m_{Y}$ $\text{Br}(\mu\to e\gamma)$ $\text{Br}(\tau\to e\gamma)$ $\text{Br}(\tau\to\mu\gamma)$ $\text{CR}(\mu^{-}\mathrm{Al}\to e^{-}\mathrm{Al})$
$4.13$ $4.29$ $4.69$ $3.82$ $1.98$ $4.87$ $41.3$ $2.72\times 10^{-19}$ $1.21\times 10^{-18}$ $3.46\times 10^{-18}$ $7.77\times 10^{-22}$
$3.25$ $2.04$ $1.91$ $3.27$ $3.49$ $4.30$ $32.5$ $1.91\times 10^{-17}$ $9.48\times 10^{-18}$ $1.73\times 10^{-18}$ $5.46\times 10^{-20}$
$2.37$ $2.51$ $4.64$ $1.08$ $5.45$ $3.18$ $23.7$ $2.41\times 10^{-16}$ $1.21\times 10^{-16}$ $2.13\times 10^{-17}$ $6.89\times 10^{-19}$
$1.56$ $2.06$ $2.35$ $5.56$ $5.15$ $4.48$ $15.6$ $1.31\times 10^{-14}$ $2.57\times 10^{-15}$ $5.89\times 10^{-15}$ $3.74\times 10^{-17}$
$3.74$ $5.45$ $2.06$ $5.77$ $2.53$ $2.37$ $37.4$ $2.46\times 10^{-18}$ $2.83\times 10^{-18}$ $1.07\times 10^{-17}$ $7.03\times 10^{-20}$
$2.81$ $2.33$ $5.47$ $1.17$ $5.78$ $2.27$ $28.1$ $4.29\times 10^{-17}$ $2.04\times 10^{-17}$ $4.10\times 10^{-18}$ $1.22\times 10^{-19}$
$2.68$ $4.52$ $2.99$ $2.47$ $1.64$ $4.09$ $26.8$ $7.25\times 10^{-19}$ $1.23\times 10^{-19}$ $4.18\times 10^{-19}$ $2.07\times 10^{-21}$
$4.19$ $3.53$ $2.84$ $4.79$ $4.19$ $1.80$ $41.9$ $4.82\times 10^{-16}$ $7.74\times 10^{-16}$ $2.76\times 10^{-15}$ $1.38\times 10^{-18}$
$3.22$ $4.87$ $2.17$ $1.18$ $5.57$ $4.12$ $38.7$ $1.42\times 10^{-17}$ $3.96\times 10^{-17}$ $1.42\times 10^{-17}$ $4.06\times 10^{-20}$
$3.85$ $3.77$ $1.93$ $5.28$ $5.39$ $3.81$ $29.5$ $7.11\times 10^{-16}$ $4.47\times 10^{-17}$ $2.24\times 10^{-16}$ $2.03\times 10^{-18}$
$4.51$ $4.36$ $2.54$ $4.09$ $5.04$ $1.33$ $31.6$ $1.68\times 10^{-14}$ $3.01\times 10^{-14}$ $1.68\times 10^{-14}$ $4.80\times 10^{-17}$
$4.55$ $3.91$ $4.35$ $4.80$ $4.40$ $1.54$ $45.5$ $2.35\times 10^{-17}$ $3.98\times 10^{-17}$ $1.38\times 10^{-16}$ $6.71\times 10^{-20}$
$3.14$ $4.92$ $5.23$ $2.46$ $3.19$ $5.50$ $31.4$ $1.53\times 10^{-16}$ $3.43\times 10^{-16}$ $2.46\times 10^{-16}$ $4.38\times 10^{-19}$
$4.27$ $2.46$ $3.30$ $4.46$ $4.89$ $3.34$ $40.6$ $5.39\times 10^{-18}$ $7.07\times 10^{-18}$ $5.35\times 10^{-18}$ $1.54\times 10^{-20}$
$2.88$ $4.77$ $5.23$ $3.95$ $1.93$ $5.14$ $28.8$ $8.31\times 10^{-19}$ $9.14\times 10^{-18}$ $3.38\times 10^{-18}$ $2.38\times 10^{-21}$
@@
Table 5: Branching ratios for the LFV decays. The second to seventh columns present the numerical values of the heavy neutrino masses $m_{n_{i}}$ in TeV. The eighth column shows the mass of the gauge boson $m_{Y}$ in TeV. The last columns display the branching ratios for the processes $\mu\to e\gamma$, $\tau\to e\gamma$, and $\tau\to\mu\gamma$, respectively. The final column shows the estimated conversion for the process $\mu^{-}\mathrm{Al}\to e^{-}\mathrm{Al}$.
VII Conclusions
@@
The $A_{4},Z_{3},Z_{4}$ and $Z_{7}$ symmetries ensure that the top quark is generated from a renormalizable Yukawa interaction, whereas the masses of lighter fermions arise from non-renormalizable Yukawa interactions after the spontaneous breaking of discrete symmetries. In addition, these discrete symmetries allow for a reduction of fermion sector parameters and yield a hierarchical structure in the entries of the SM charged fermion mass matrices, which effectively explains the SM charged fermion masses and quark mixing pattern. Due to the discrete symmetries of the model, the quark mixing arises only from the up-type quark sector. The $Z_{10}$ symmetry distinguishes the $A_{4}$ scalar triplets participating in the quark and neutrino Yukawa interactions from the ones appearing in the charged lepton Yukawa terms, then allowing to separate the mixing in the neutrino and quark sectors. The model successfully predicts SM fermion masses and mixing patterns, consistent with experimental observations. The sizable neutrino Yukawa couplings predicted by the model could lead to observable rates of cLFV processes, which could be tested in future experiments.
@@ Acknowledgments
The authors are very grateful to L.T. Hue, for his involvement in the initial stages of the project and for very useful discussions. This research has received funding from Fondecyt (Chile), Grants No. 1210378, No. 1241855, No 1210131, ANID PIA/APOYO AFB230003 and Proyecto Milenio-ANID: ICN2019_044. D.T.Huong acknowledges the financial support of the Vietnam Academy of Science and Technology under Grant No. NVCC05.05/24-25; No. ICP-2025.02.

### v2

@@ VII Lepton flavor violating decays $e_{j}\rightarrow e_{i}$
Figure 8: Correlation between the branching ratios $\text{Br}(\mu\to e\gamma)$ and $\text{Br}(\tau\to e\gamma)$ (panel (a)), $\text{Br}(\tau\to\mu\gamma)$ and the conversion rate $\mathrm{CR}(\mu^{-}\text{Al}\to e^{-}\text{Al})$ (panel (b)) and $\text{Br}(\tau\to e\gamma)$ and the conversion rate $\mathrm{CR}(\mu^{-}\text{Al}\to e^{-}\text{Al})$. The color gradient represents the values of $\text{Br}(\tau\to\mu\gamma)$ in panel (a), the gauge boson mass $m_{Y}$ in panel (b) and the charged scalar mass $m_{\rho_{3}^{\pm}}$ in panel (c). The yellow shaded region corresponds to the exclusion zone from the MEG II experiment [142], the red dot dashed line indicates the current bound for $\mu\to e$ conversion in titanium (Ti) nuclei [149] and the gray dashed line represents the boundary approximated by Eq. (118).
$m_{n_{4}}$ $m_{n_{5}}$ $m_{n_{6}}$ $m_{n_{7}}$ $m_{n_{8}}$ $m_{n_{9}}$ $m_{Y}$ $m_{\rho_{3}^{\pm}}$ $\text{Br}(\mu\to e\gamma)$ $\text{Br}(\tau\to e\gamma)$ $\text{Br}(\tau\to\mu\gamma)$ $\mathrm{CR}(\mu\mathrm{Al}\to e\mathrm{Al})$
$4.90$ $3.52$ $2.81$ $3.98$ $3.23$ $5.57$ $10.21$ $118.99$ $7.14\times 10^{-13}$ $1.79\times 10^{-13}$ $1.23\times 10^{-12}$ $2.04\times 10^{-15}$
$5.37$ $0.95$ $1.79$ $3.23$ $4.63$ $1.43$ $12.84$ $65.78$ $1.41\times 10^{-13}$ $1.16\times 10^{-13}$ $2.16\times 10^{-14}$ $4.03\times 10^{-16}$
$3.35$ $2.55$ $3.07$ $2.95$ $1.65$ $5.82$ $13.28$ $167.09$ $1.65\times 10^{-13}$ $3.57\times 10^{-14}$ $6.21\times 10^{-14}$ $4.72\times 10^{-16}$
$2.14$ $4.31$ $2.74$ $5.83$ $4.47$ $5.48$ $10.56$ $173.46$ $3.16\times 10^{-14}$ $1.96\times 10^{-13}$ $3.12\times 10^{-13}$ $9.04\times 10^{-17}$
$1.47$ $2.34$ $1.64$ $4.60$ $1.78$ $4.42$ $11.22$ $150.23$ $2.59\times 10^{-14}$ $4.71\times 10^{-14}$ $1.61\times 10^{-13}$ $7.41\times 10^{-17}$
$1.01$ $5.78$ $4.70$ $4.29$ $1.95$ $5.84$ $17.61$ $118.51$ $9.32\times 10^{-15}$ $7.66\times 10^{-15}$ $1.44\times 10^{-15}$ $2.66\times 10^{-17}$
$4.71$ $3.30$ $2.58$ $3.73$ $1.70$ $3.47$ $14.00$ $177.13$ $7.98\times 10^{-15}$ $2.41\times 10^{-14}$ $2.17\times 10^{-14}$ $2.28\times 10^{-17}$
$2.94$ $3.44$ $1.41$ $4.70$ $4.79$ $5.47$ $15.59$ $39.34$ $3.17\times 10^{-16}$ $4.58\times 10^{-16}$ $1.95\times 10^{-16}$ $9.05\times 10^{-19}$
$2.40$ $1.45$ $5.88$ $5.92$ $4.47$ $2.16$ $18.42$ $28.29$ $6.02\times 10^{-16}$ $8.22\times 10^{-16}$ $3.11\times 10^{-15}$ $1.72\times 10^{-18}$
$4.33$ $4.22$ $4.41$ $4.93$ $2.57$ $3.17$ $25.87$ $63.28$ $1.86\times 10^{-17}$ $7.04\times 10^{-17}$ $7.33\times 10^{-17}$ $5.30\times 10^{-20}$
$1.21$ $2.22$ $1.32$ $4.38$ $1.35$ $3.98$ $23.03$ $42.50$ $1.98\times 10^{-17}$ $5.38\times 10^{-17}$ $1.46\times 10^{-16}$ $5.65\times 10^{-20}$
$3.19$ $1.92$ $1.15$ $3.89$ $3.59$ $1.22$ $37.17$ $126.11$ $5.27\times 10^{-18}$ $3.43\times 10^{-18}$ $5.48\times 10^{-19}$ $1.51\times 10^{-20}$
$2.43$ $2.85$ $5.25$ $0.95$ $3.76$ $3.65$ $45.50$ $34.87$ $3.27\times 10^{-18}$ $1.97\times 10^{-18}$ $5.58\times 10^{-19}$ $9.36\times 10^{-21}$
$4.75$ $2.02$ $1.03$ $4.09$ $4.58$ $4.68$ $45.53$ $55.09$ $3.63\times 10^{-19}$ $6.12\times 10^{-19}$ $3.25\times 10^{-19}$ $1.04\times 10^{-21}$
$3.99$ $1.85$ $3.44$ $3.75$ $5.96$ $4.04$ $40.51$ $102.52$ $3.30\times 10^{-19}$ $1.87\times 10^{-19}$ $9.21\times 10^{-19}$ $9.41\times 10^{-22}$
@@
Table 5: Branching ratios for the LFV decays. The first to sixth columns present the numerical values of the heavy neutrino masses $m_{n_{i}}$ in TeV. The seventh and eighth columns show the mass of the gauge boson $m_{Y}$ and the charged scalar mass $m_{\rho_{3}^{\pm}}$ in TeV, respectively. The subsequent columns display the branching ratios for the processes $\mu\to e\gamma$, $\tau\to e\gamma$, and $\tau\to\mu\gamma$, respectively. The final column shows the estimated conversion for the process $\mu\mathrm{Al}\to e\mathrm{Al}$.
In what follows we provide a qualitative discussion about the implications of our model in the muon anomalous magnetic moment. It is important to note that the muon anomalous magnetic moment receives several contributions. These include effects from the virtual exchange of heavy neutral and electrically charged gauge bosons paired with their corresponding charged and neutral leptons, as well as contributions from electrically charged scalars and neutrinos. However, these additional contributions are negligible due to the highly suppressed mixing angle between the neutral CP-even component of the $SU(3)_{L}$ scalar triplet $\rho$ and the other CP-even scalars. As for the contribution from electrically charged scalars and light active neutrinos, our numerical analysis shows that it can account for the observed magnitude of the muon $g-2$ anomaly when the charged scalars are lighter than $400$ GeV. Nevertheless, this contribution is negative and therefore cannot explain the correct sign of the anomaly. In order to successfully reproduce the $g-2$ muon anomaly in our model, its fermion sector has to be enlarged by the inclusion of charged vector like leptons, transforming as singlets under the $SU(3)_{C}\times SU(3)_{L}$ gauge symmetry, in a similar way as done in the 331 model with $D_{4}$ flavor symmetry of Ref. [150]. Those charged vector like leptons should have the appropiate transformation properties under the discrete groups of our model in such a way to allow them having Yukawa interactions with the SM charged leptons. This will give rise to one loop level contributions to the muon anomalous magnetic moment mediated by charged vector like leptons and neutral scalars running in the internal lines of the loop. Under these conditions the muon $g-2$ anomaly can be successfully accommodated for an appropiate region of parameter space provided that such exotic charged vector like leptons are included in the fermionic spectrum of our model. We do not perform such fermionic extension in this work since recent lattice QCD calculations of the hadronic vacuum polarization (HVP) contribution of the muon magnetic moment [151, 152, 153, 154] lead to a $1.5\sigma$ deviation [155] of its SM prediction from the corresponding experimental value and in addition, a recent study of the hadronic vacuum polarization contribution to the muon $g-2$ at long distances has found that the SM prediction for the total muon anomalous magnetic moment is in agreement with the current experimental average [156].
Although the model includes sources of CP violation through complex Yukawa couplings and active–sterile neutrino mixing, the contributions to the electron electric dipole moment (EDM) are strongly suppressed. This suppression is due to the large masses of the new charged scalar $\rho_{3}^{\pm}$ and the exotic gauge boson $Y^{\pm}$, which, due to flavour constraints, are typically required to lie above the multi-TeV scale. As pointed out in Ref. [157], the corresponding Wilson coefficients scale inversely with the square of these masses. Furthermore, the structure imposed by the $A_{4}\times Z_{N}$ flavour symmetries tends to align the Yukawa couplings and reduce the number of physical CP-violating phases. Together with the small active–sterile mixing parameters resulting from the bounds imposed by charged lepton flavour-violating processes, it follows that the predicted contributions to the electric dipole moment $d_{e}$ of the electron are well below the current experimental sensitivity. Therefore, the electron EDM does not place any significant constraint on the model parameter space.
To conclude this section, we present some remarks on the decays of quasi-Dirac sterile neutrinos and compare our predictions for these decays and lepton flavor violating (LFV) signals with those obtained in other models featuring extended gauge symmetries. Similar to the $U^{\prime}(1)$ model discussed in Ref. [158], our model predicts that sterile neutrinos undergo two-body decays: $N\rightarrow l^{\pm}_{i}W^{\mp}$, $\nu_{i}Z$ and $\nu_{i}h$ (where $i=1,2,3$ is a flavor index). These decay modes are suppressed due to the small active-sterile neutrino mixing angle, $\theta\sim\mathcal{O}(10^{-2})$. The above mentioned two body decays lead to three-body decays: $N\rightarrow l^{+}_{i}l^{-}_{j}\nu_{k}$, $N\rightarrow l^{-}_{i}u_{j}\bar{d}_{k}$, $N\rightarrow b\bar{b}\nu_{k}$ (where $i,j,k=1,2,3$ are flavor indices), which are also present in the $U^{\prime}(1)$ model of Ref. [158]. As a result, we anticipate similar predictions for the total cross section of the LFV process $pp\rightarrow NN\rightarrow e^{\pm}\mu^{\mp}4j$, as well as for sterile neutrino decays, compared to those found in Ref. [158]. However, a slight deviation is expected in the decay rate of $N\rightarrow l^{+}_{i}l^{-}_{j}\nu_{k}$, as our model includes contributions from off-shell $W$ and $W^{\prime}$ gauge bosons whereas in Ref. [158], only the off-shell $W$ boson contributes. Notably, the contribution from the off-shell $W^{\prime}$ boson in our model is significantly suppressed by a factor of approximately $\frac{M^{4}_{W}}{M^{4}_{W^{\prime}}}$ relative to the contribution from the $W$ boson. Furthermore, we anticipate similar predictions for sterile neutrino decay rates and the cross section of the LFV process $pp\rightarrow NN\rightarrow e^{\pm}\mu^{\mp}$ to those found in the left-right symmetric model of Refs. [126, 127]. However, the charged lepton flavor violating process $\mu\to e\gamma$ serves as a key discriminator between our model and the left-right symmetric model of Refs. [126, 127]. In our case, the dominant contributions to $\mu\to e\gamma$ originates from loop diagrams involving $W^{\prime}$ and charged scalar exchanges, whereas in the left-right symmetric models of Refs. [126, 127], such CLFV process receives contributions from the virtual exchange of the $W_{R}$ gauge boson and the doubly charged scalars running in the internal lines of the loop.
VIII Conclusions
@@
The $A_{4},Z_{3},Z_{4}$ and $Z_{7}$ symmetries ensure that the top quark is generated from a renormalizable Yukawa interaction, whereas the masses of lighter fermions arise from non-renormalizable Yukawa interactions after the spontaneous breaking of discrete symmetries. In addition, these discrete symmetries allow for a reduction of fermion sector parameters and yield a hierarchical structure in the entries of the SM charged fermion mass matrices, which effectively explains the SM charged fermion masses and quark mixing pattern. Due to the discrete symmetries of the model, the quark mixing arises only from the up-type quark sector. The $Z_{10}$ symmetry distinguishes the $A_{4}$ scalar triplets participating in the quark and neutrino Yukawa interactions from the ones appearing in the charged lepton Yukawa terms, then allowing us to separate the mixing in the neutrino and quark sectors. The model successfully predicts SM fermion masses and mixing patterns, consistent with experimental observations. The sizable neutrino Yukawa couplings predicted by the model could lead to observable rates of cLFV processes, which could be tested in future experiments.
@@ Acknowledgments
The authors are very grateful to L.T. Hue, for his involvement in the initial stages of the project and for very useful discussions. This research has received funding from Fondecyt (Chile), Grants No. 1210378, No. 1241855, No 1210131, ANID PIA/APOYO AFB230003 and Proyecto Milenio-ANID: ICN2019_044. D.T.Huong acknowledges the financial support of the Vietnam Academy of Science and Technology under Grant No. NVCC05.05/24-25.